# LERa: Replanning with Visual Feedback in Instruction Following

### v1

@@ IV-B Metrics
We use the following metrics to evaluate the replanning quality: Success Rate (SR), Goal-Condition Success (GCR), and Success Rate of Replanning (SRep). SR is the ratio of the number of successfully completed episodes to the total number of episodes. An episode is considered successfully completed when all target objects are in their target positions and states at the end of the episode. For ALFRED and VirtualHome this mean all objects are in correct places and in specific state, e.g. apple is cutted and fridge is closed. For PyBullet target states are correct position of blocks in bowls, e.g. blue block on red one and both in yellow bowl. GCS is the ratio of the number of goal conditions successfully met at the end of an episode to all goal conditions in that episode, averaged over the number of episodes. This metric can be interpreted as fraction of all objects that met their target state. SRep is the ratio of the number of successful replannings module execution to the number of all replannings in an episode, averaged over the number of episodes. A replanning is considered successful if it results in the agent continuing to execute the plan resolving issue.
@@ IV-C Agents
The LERa approach functions solely as a replanning module, meaning it cannot be used for initial plan generation with guarantees on its exutability. To evaluate the replanning capability of LERa, we incorporate ground-truth plans with agents that achieve the highest success rates in their respective environments.
In our experiments across all environments, we employed an Oracle agent as the main baseline. This agent is provided with a complete initial plan that solves the task assuming that there are no objects with altered states or execution errors. Those plans are predefined and known for each task in all environments.
However, the agent not able to replan and strictly follows the predefined plan. If an action in the plan cannot be executed, the agent simply moves on to the next action. In the VirtualHome and PyBullet environments, the agent achieves a 100% success rate (SR) as long as there are no changes in the scene. In the ALFRED environment, when using ground-truth (GT) plans, the agent successfully completes tasks with an SR of approximately 75%. Although we use the GT plans, ALFRED’s predefined templates are suboptimal, as they do not account for hidden objects in receptacles. For example, if an agent is retrieving a knife from one of multiple drawers, the oracle may teleport to an empty one. These imperfections reduce SR by approximately 15–20%. An additional $\approx 5\%$ drop results from AI2THOR bugs, such as unexplained “PutObject” failures or objects being rendered without corresponding masks.
@@ IV-D Baselines
To assess the contribution of each step in LERa, we conducted experiments with its different variations – Ra, LRa, and ERa – each representing a distinct combination of replanning steps. These experiments allow for a detailed analysis of the role of each component and provides a structured comparison against the full method. Each LERa variation is an exact replica of the corresponding part of the main approach, ensuring consistency in evaluation. The ERa variant, in particular, represents one of the most commonly used error correction strategies. It does not process visual information but instead relies on textual feedback from the environment, which explicitly indicates whether a given action was unsuccessful. This makes ER a widely applicable baseline, as it simulates a scenario which the agent corrects errors based solely on system-provided failure signals rather than direct perception. The Baseline represents a commonly used method in recent research [12, 14]. It is not a direct implementation of any specific approach, but rather an abstraction of the one-shot plan prediction paradigm, where the VLM receives all available information in a single step.
It performs immediate replanning upon receiving visual information and additional task-related data, directly outputting a corrected plan. In our implementation, it corresponds to the “Replan” step of LERa, with the addition of visual feedback. However, unlike LRa, the baseline operates in a single-step manner, generating a new without intermediate reasoning.
@@ IV-E Ablations
In the additional experiments, we demonstrate how the LERa module performs when other modules make mistakes. In real world scenarious it is almoust impossible to create checker detects errors with absolute accuracy. Thus The first three agents (“O-05”, “O-10” and “O-15”) have imperfect Subtask Checker that flips the prediction with probabilities of 0.05, 0.10 and 0.15, respectively. The forth agent (“O-FC”) uses FILM’s [21] Subtask Checker that has check error probability of $\approx 2.5\%$.
@@ IV-F Implementation Details
The experiments were run on two NVIDIA RTX Titan 26 Gb. One GPU was running a VLM and the other – a virtual environment and an ALFRED agent. MiniGPT-v2 [22], was used for experiment with imperfect checkers. We chose this model due to its high image understanding abilities and small size. The generation hyperparameters were set so that the responses were deterministic. Virtual Home and PyBullet does not require much computational resources and was running on local PC. For experiments with proprietary models, their public API was used. For experiments we use gpt-4o and gpt4o-mini [23], Gemini-Flash-1.5 and Gemini-Pro-1.5 [24], LLaMA-3.2 [25].
@@ V-A LERa Replanning Abilities
Our experiments demonstrate that the LERa approach significantly improves task solving success rates across all scenarios. Table I presents results for the Oracle, Baseline, LERa and its variations. In VirtualHome-ChaOS LERa successfully solves almost all tasks with 94% SR. ALFRED-ChaOS, being more complex, in terms of plan structure, environment than VirtualHome, have a more moderate improvement; however, with LERa Oracle agent solves 50% more tasks. In PyBullet, where single action may fail even after replanning due to stochasticity the impact of replanning is the most significant.
Experiments results with LERa variations – Ra, ERa, and LRa – demonstrate that most tasks can be solved without the “Explain” step in PyBullet environments. The main reason is simplicity of PyBullet action. A detailed analysis is provided in Section V-B. We were surprised to find that LR successfully solves tasks in PyBullet. To verify the consistency of these results, we conducted additional experiments using the Gemini model. The results from the PyBullet-Gemini experiments consistently indicates that replanning has a significant impact on task success rates. LERa outperforms both Oracle and the baseline agents. This leads to conclusion that leveraging most suitable state-of-the-art language models enhances our approach effectiveness across diverse domains.
@@ V-B Visual Feedback Impact
Experiments in ALFRED, as our first virtual environment, confirmed that visual information is critical for error understanding and further correction. However, without structured output formatting and explicit failure reasoning, tasks remain unsolvable. The O-ERa agent fails to handle simple tasks due to an incomplete understanding of the environment, while O-LRa generates clear failure explanations but lacks stable and reasonable planning. Combining these approaches led to the LERa module and the O-LERa agent, achieving the best results. The three-step replanning process, which included extracting error-relevant information, reasoning about the failure, and performing format constrained replanning, proved effective. Later experiments in VirtualHome confirmed these findings. PyBullet is more straightforward environment with less available actions and short object list. The O-LRa achieved the best results, finally predicting stable and well formated action plans. Despite this, the three-step approach remains preferable due to its stability, interpretability, and consistency.
Figure 6: Hard replanning scenario in TableTop-RoboticStand experiments. Steps 2 and 4 mark error occurrences. In step 3, LERa intentionally repeats the previous action, attempting a successful retry. In step 5, after the pear slightly moved, the module successfully identified it as not being an apple. Consequently, LERa excluded the “pick apple” action from the plan, as no apple was present on the table, and proceeded to execute the subsequent steps.
@@ V-C Impact of Different VLMs
The results of experiments with different VLMs are presented in Table II. Although LERa does not require training and operates using off-the-shelf VLMs, its performance is highly dependent on the accuracy of the VLM’s image analysis. A consistent decline in performance is observed across all environments as the quality of the VLM picture analysis decreases. Additionally, the choice of VLM affects performance depending on the complexity of the environment. For simpler environments, such as PyBullet, Gemini models demonstrate superior results. In contrast, for more complex environments requiring advanced reasoning, GPT-based models perform better.
Although LERa involves multiple model queries, replanning is not required at every step, keeping overall system load low. Latency can be further reduced through parallelism in future implementations.
@@ V-D Impact of Non-Perfect Checker
In real-world applications, designing a flawless action execution checker is challenging. In our series of additional experiments, we demonstrate how the LERa module performs when other modules make mistakes. In Table III it is easy to see that the use of the module significantly improves the success rate in all cases. Despite the high error rate of Subtask Checker, LERa is able to cope with various situations that were not described in few-shot prompts.
@@ V-E Results on the Real Robot
To evaluate the performance of the LERa module, we conducted a series of experiments on a robotic arm tasked with manipulating various objects. The purpose of these experiments was to test the capabilities of the replanning in scenarios with unexpected errors. A total of 18 series were conducted, each of which included different configurations of objects and targets. The agent successfully completed the renovation in 15 of the 18 series, demonstrating its reliability and adaptability in dynamic conditions. Fig. 6 illustrates an example of a plan execution. The initial plan specified the task: “remove the apple and the lemon from the table”. The challenge in this scenario arises from a misclassification made by the detection module, which identified a green, round object – a pear – as an apple, incorporating it into the initial plan. Since the manipulator was unable to lift the pear, execution errors occurred. After the pear slightly rotated, the LERa module re-evaluated the object and correctly determined it was not an apple, subsequently omitting the steps in the plan related to the misclassified object.
@@ V-F Failure Analysis
In ALFRED, one of the most common and significant challenges in replanning is the inability to infer the cause of failure from observations. Unlike in VirtualHome, where the agent’s camera is positioned slightly behind, providing a clear view of objects in most cases, ALFRED employs a first-person perspective. This often leads to ambiguity in perception, making it difficult for the replanner to determine whether the observed scene includes an open refrigerator, a microwave, or simply an empty shelf. The VirtualHome failure case arose from similar problems, with the VLM misinterpreting the opened microwave door as a closed transparent one, likely due to its low position partially excluding the door from the frame. In PyBullet, most errors arise when the VLM fails to predict the necessary “locate object” action before attempting to pick or place it, or when it does not fully resolve the initial instruction. The model may generate a new plan that corrects an observed dropping error but fails to account for objects already placed in their correct positions. As a result, the agent incorrectly relocates blocks that were previously positioned correctly. This behavior was particularly noticeable in the outputs of smaller models, whereas GPT and Gemini demonstrated a better understanding of why already placed blocks should not be moved. Additionally, despite all random parameters being fixed, some intrinsic environment stochasticity was observed. In certain cases, blocks that were randomly dropped fell into positions where the gripper was unable to pick them up, leading to further task failures.
@@ VI CONCLUSION
Large Language Models excel in robotic task planning but struggle with dynamic changes and execution failures due to their reliance on textual input. To address this, we introduce LERa – a Visual Language Model-based replanning approach that leverages visual feedback for effective replanning. Unlike existing methods, LERa requires minimal additional information, making it a more practical and adaptable solution across different agent architectures. We also propose three new environments, including ALFRED-ChaOS and VirtualHome-ChaOS, to evaluate replanning in dynamic settings. Extensive experiments show that LERa outperforms baselines in handling environmental and execution errors. Additionally, we integrate LERa into a tabletop manipulation robot, validating its effectiveness in real-world scenarios. Our findings highlight the robustness and generalizability of LERa. In future work, we will extend its capabilities to handle a broader range of errors, especially in long-horizon tasks requiring complex temporal and spatial reasoning. We believe that our work not only advances VLM-based replanning but also provides valuable benchmarks for future research in robotic task execution.

### v2

@@ IV-B Metrics
We use the following metrics to evaluate replanning quality: Success Rate (SR), Goal-Condition Success (GCR), and Success Rate of Replanning (SRep). SR is the ratio of the number of successfully completed episodes to the total number of episodes. An episode is considered successfully completed when all target objects are in their target positions and states at the end of the episode. For ALFRED and VirtualHome, this means all objects are in the correct places and in the specific states, e.g., the apple is cut and it is in the closed fridge. For PyBullet, the target states are the correct positions of blocks in bowls, e.g., the blue block is on the red one and both are in the yellow bowl. GCS is the ratio of the number of goal conditions successfully met at the end of an episode to all goal conditions in that episode, averaged over the number of episodes. This metric can be interpreted as the fraction of all objects that met their target state. SRep is the ratio of the number of successful replanning attempts to the total number of replanning attempts in an episode, averaged over the number of episodes. Replanning is considered successful if it results in the agent continuing to execute the plan and resolving the issue.
@@ IV-C Agents
The LERa approach functions solely as a replanning module, meaning it cannot be used for initial plan generation with guarantees of effectiveness. To evaluate the replanning capability of LERa and isolate its replanning performance from errors in other modules, we used ground-truth plans and oracle agents. Oracle agents are those that have perfect or near-perfect knowledge of the environment and execute actions with the maximum achievable probability of success (usually 100%).
We used oracle agents for each virtual environment in our experiments and the agents that were not able to replan are denoted as “Oracle”. These agents were provided with a complete, predefined initial plan that solves the task assuming there are no objects with altered states or execution errors. Therefore, if an action in the plan cannot be executed, the Oracle agents simply moved on to the next action, eventually leading to a failed episode. In the VirtualHome and PyBullet environments, the agents achieved a 100% success rate (SR) as long as there were no changes in the scenes. In the ALFRED environment, the Oracle agent successfully completed tasks with an SR of approximately 75%. Although the GT plans were used, ALFRED’s predefined templates are suboptimal, as they do not account for objects hidden in receptacles. For example, if the task is to retrieve a knife from one of multiple drawers, Oracle may teleport to an empty one. These imperfections reduced SR by approximately 15–20%. An additional $\approx 5\%$ drop resulted from AI2THOR bugs, such as unexplained “PutObject” failures or objects being rendered without corresponding masks.
@@ IV-D Baselines
To assess the contribution of each step in LERa, we conducted experiments with its different variations – Ra, LRa, and ERa – each representing a distinct combination of replanning steps. These experiments allow for a detailed analysis of the role of each component and provide a structured comparison against the full method. Each LERa variation is an exact replica of the corresponding part of the main approach, ensuring consistency in evaluation. The ERa variant, in particular, represents one of the most commonly used error correction strategies. It does not process visual information but instead relies on textual feedback from the environment that explicitly indicates whether a given action was unsuccessful. This makes ERa a widely applicable baseline as it simulates a scenario in which the agent corrects errors based solely on system-provided failure signals rather than on direct perception.
Baseline represents a method commonly used in recent research [12, 14]. It is not a direct implementation of any specific approach, but rather an abstraction of the one-shot plan prediction paradigm in which the VLM receives all available information in a single step.
The method immediately replans upon receiving visual information and additional task-related data, directly outputting a corrected plan. In our implementation, it corresponds to the “Replan” step of LERa, with the addition of visual feedback. However, unlike LRa, Baseline operates in a single-step manner, generating a new plan without intermediate reasoning.
@@ IV-E Ablations
In the additional experiments, we demonstrate how the LERa module performs when other modules make mistakes. In real-world scenarios, it is nearly impossible to create a checker that detects errors with absolute accuracy. Thus, the first three agents (“O-05”, “O-10” and “O-15”) have the imperfect Task Checker that flips the true prediction with probabilities of 0.05, 0.10 and 0.15, respectively. The fourth agent (“O-FC”) uses FILM’s [21] Task Checker that has a check error probability of $\approx 2.5\%$.
@@ IV-F Implementation Details
We used public APIs for experiments with proprietary models (gpt-4o and gpt4o-mini [22], Gemini-Flash-1.5 and Gemini-Pro-1.5 [23]) and two NVIDIA RTX Titan 26 GB for experiments with open-source models (MiniGPT-v2 [24], LLaMA-3.2 [25]): one GPU handled a VLM, and the other managed a virtual environment and an agent. MiniGPT-v2 was used in experiments with imperfect checkers. The generation hyperparameters were configured to ensure deterministic responses. VirtualHome and PyBullet do not require significant computational resources and were therefore run on a local PC.
@@ V-A LERa Replanning Abilities
Our experiments demonstrate that the LERa approach significantly improves task-solving success rates across all scenarios over Baseline. Table I presents the results for Oracle, Baseline, LERa, and its variations. In VirtualHome-ChaOS, LERa successfully solves almost all tasks, achieving 94% SR. ALFRED-ChaOS, being a more complex environment than VirtualHome in terms of plan structure, has a more moderate improvement; however, the Oracle agent solves 50% more tasks with LERa. In PyBullet, where a single action may fail even after replanning due to stochasticity, the impact of replanning is the most significant.
Experiments results with LERa variations — Ra, ERa, and LRa — demonstrate that most tasks can be solved without the “Explain” step in PyBullet scenarios. The main reason is the simplicity of PyBullet’s actions. A detailed analysis is provided in Section V-B. We were surprised to find that LRa solves tasks better than LERa in PyBullet. To verify the consistency of these results, we conducted additional experiments using the Gemini model. The results from the PyBullet-Gemini experiments consistently indicate that replanning has a significant impact on task success rates.
@@ V-B Visual Feedback Impact
Experiments in ALFRED confirmed that visual information is critical for understanding and correcting errors. However, tasks remain unsolvable without explicit failure reasoning and structured output formatting. The O-ERa agent fails to handle simple tasks due to an incomplete understanding of the environment, while O-LRa generates clear failure explanations but lacks stable and reasonable planning. Combining these approaches led to the LERa module and the O-LERa agent, which achieved the best results. The three-step replanning process, which includes extracting error-relevant information, reasoning about the failure, and performing format-constrained replanning, proved effective. Later experiments in VirtualHome confirmed these findings.
PyBullet is a more straightforward environment with fewer available actions and a shorter object list. The O-LRa achieved the best results, finally predicting stable and well-formatted action plans. Despite this, the three-step approach remains preferable due to its consistency and interpretability.
Figure 6: Hard replanning scenario in TableTop Robotic Stand experiments. Steps 2 and 4 mark error occurrences. In step 3, LERa intentionally repeats the previous action, attempting a successful retry. In step 5, after the pear moved slightly, the module successfully identified it as not an apple. Consequently, LERa excluded the “pick apple” action from the plan, as no apple was present on the table, and proceeded to execute the subsequent steps.
@@ V-C Impact of Different VLMs
The results of experiments with different VLMs are presented in Table II. Although LERa does not require training and operates using off-the-shelf VLMs, its performance is highly dependent on the accuracy of the VLM’s image analysis. A consistent decline in performance is observed across all environments as the quality of the VLM’s picture analysis decreases. Additionally, the choice of the VLM affects performance depending on the complexity of the environment. For simpler environments, such as PyBullet, Gemini models demonstrate superior results. In contrast, GPT-based models perform better in more complex environments that require advanced reasoning.
Although LERa involves multiple model queries, replanning is not required at every step, keeping the overall system load low. Latency can be further reduced through model inference optimization in future implementations.
@@ V-D Impact of Non-Perfect Checker
In real-world applications, designing a flawless action execution checker is challenging. In our series of additional experiments, we demonstrate how the LERa module performs when other modules make mistakes. Table III shows that the use of the module significantly improves the success rate in all cases. Despite errors of the task checkers, LERa is able to cope with various situations not described in the few-shot prompts.
@@ V-E Results on the Real Robot
To evaluate the performance of the LERa module, we conducted a series of experiments using the robotic arm tasked with manipulating various objects. The purpose of these experiments was to test replanning capabilities in scenarios with unexpected errors. A total of 18 series were conducted, each including different configurations of objects and targets. The agent successfully completed the replanning in 15 of the 18 series, demonstrating its reliability and adaptability under dynamic conditions. Fig. 6 illustrates an example of a plan execution. The initial plan specified the task: “Remove the apple and the lemon from the table”. The challenge in this scenario arose from a misclassification made by the detection module, which identified a green, round object — a pear — as an apple and incorporated it into the initial plan. Since the manipulator was unable to lift the pear, execution errors occurred. After the pear slightly rotated, the LERa module re-evaluated the object and correctly determined that it was not an apple, subsequently omitting the steps in the plan related to the misclassified object.
@@ V-F Failure Analysis
In ALFRED, one of the most common and significant challenges in replanning is the inability to infer the cause of failure from observations. Unlike in VirtualHome, where the agent’s camera is positioned slightly behind, providing a clear view of objects in most cases, ALFRED employs a first-person perspective. This often leads to ambiguity in perception, making it difficult for the replanner to determine whether the observed scene includes an open refrigerator, a microwave, or simply an empty shelf. The VirtualHome failure cases arose from similar problems, with the VLM misinterpreting the opened microwave door as a closed transparent one, likely due to its low position partially excluding the door from the frame. In PyBullet, most errors occurred when the VLM failed to predict the necessary “locate object” action before attempting to pick or place it, or when it did not fully resolve the initial instruction. The model could generate a new plan that corrected an observed drop error but failed to account for objects already placed in their correct positions. As a result, the agent incorrectly relocated blocks that were previously positioned correctly. This behavior was particularly noticeable in the outputs of smaller models, whereas GPT and Gemini demonstrated better understanding of why already placed blocks should not be moved. Additionally, despite all random parameters being fixed, some intrinsic environmental stochasticity was observed. In certain cases, blocks that were randomly dropped fell into positions where the gripper was unable to pick them up, leading to further task failures.
@@ VI CONCLUSION
Large Language Models excel in robotic task planning but struggle with dynamic changes and execution failures due to their reliance on textual input. To address this, we introduce LERa — a Visual Language Model-based replanning approach that leverages visual feedback for effective replanning. Unlike existing methods, LERa requires minimal additional information, making it a more practical and adaptable solution across different agent architectures. We also propose three new environments, including ALFRED-ChaOS and VirtualHome-ChaOS, to evaluate replanning in dynamic settings. Extensive experiments show that LERa outperforms baselines in handling environmental and execution errors. Additionally, we integrate LERa into a tabletop manipulation robot, validating its effectiveness in real-world scenarios. Our findings highlight the robustness and generalizability of LERa. In future work, we will extend its capabilities to handle a broader range of errors, especially in long-horizon tasks requiring complex temporal and spatial reasoning. We believe that our work not only advances VLM-based replanning but also provides valuable benchmarks for future research in robotic task execution.
ACKNOWLEDGMENTS
The study was supported by the Ministry of Economic Development of the Russian Federation (agreement with MIPT No. 139-15-2025-013, dated June 20, 2025, IGK  [^00] C313925P4B0002).